OpenStudy (anonymous):

(5t^2+t over t-1)^2 over means fraction... find the power @Calculus1

OpenStudy (agreene):

\[(\frac{5t^2+t}{t-1})^2=\frac{(5t^2+t)^2}{(t-1)^2}\] \[=\frac{25t^4+10t^3+t^2}{t^2-2t+1}\] if you need to further expand, you can break this into its parts (break denominator) and afterwards you should arrive at: \[25 t^2 + \frac{36}{(-1 + t)^2} + \frac{132}{(-1 + t)} + 60 t +96 \]

OpenStudy (anonymous):

@agreene thanks youre a life saver!